George Q. Farmer grows soybeans. He estimates his October harvest at 55,000 bushels. The spot price of soybeans is currently $9.23/bushel. Soybean futures are defined as 5,000bu, $0.01/bu. November soybeans are quoted at 922. To fully hedge his position George should

Buy / Sell ?

How many contracts?  

The basis is $ ?

A textbook Hedge will lock in an effective price of $ ?
Answer #1

The farmer estimates that his crop will be 55000 bushels . As the farmer is long on the asset in future, to hedge his position, he has to short i.e. SELL soyabean futures,

The no. of contracts = size of crop/size of one futures contract = 55000/5000 = 11 ontracts

Basis is defined as Basis= Spot price -Futures price

  Now, Futures price = 922 (in $0.01/bushel)

=$9.22/bushel

Therefore, Basis = $9.23 -$9.22  

=$ 0.01/bushel

For one contract, the basis is $0.01/bushel * 5000 bushel = $50

The total Basis for the entire 11 contracts is 55000 bushel *$0.01/bushel = $550  

A textbook hedge will effectively lock in today's future price i.e. $9.22/bushel